The rise of olobalization is due to the many companies that have become multinational corporations for various reasons-for example, to access betber technology, to enter new markets, to obtain more raw materlals, to find funding resources, to minimize production costs, or to diverssity business risk This multimarket presence exposes companies to dfferent kinds of risk as weli-for example, political risk and exchange rate risk The relationship between interest rates and exchange rates can be represented through the concept of interest rate parity, Consider the following: Suppose you observe the following spot and forward exchange rates between the U.S. dolar ($) and the Canadian dollar (CS) One-Year Forward Spot Exchange Rate Exchange Rate Canadian dollar (U.S. dollar/Canadian dollar) 0.9001 0.8842 Ha The current one-year interest rate on U.S. Treasury securities is 8.03 %. If interest rate parity hoids, what is the expected yield on one-year Canadian securities of equal risk? Mai O 6.12 % 6.43 % om S.81 % O5.51 % Interest rate parity recognizes that when you invest in a country other than your home country, two factors affect your investment-returns on the hsgn rate Which of the following would cause the overall return on your investment to be lower than the
5.51% Interest rate parity recognizes that when you invest in a country other than your home country, two factors affect your investment-returns on the investment itself and changes in the exchange rate. Which of the following would cause the overall return on your investment to be lower than the investment's stated return? Your home currency depreciates relative to the currency in which the investment is denominated The currency in which the investment is denominated appreciates relative to your home currency. veur home currency appreciates relative to the currency in which the investment is denominated.
The rise of globalization is due to the many companies that have become multinational corporations for various reasons-for example, to access better technology, to enter new markets, to obtain more raw materials, to find funding resources, to minimize production costs, or to diversify business risk This multimarket presence exposes companies to different kinds of risk as well-for example, political risk and exchange rate risk The relationship between interest rates and exchange rates can be represented through the concept of interest rate parity. Consider the following: Suppose you observe the following spot and forward exchange rates between the U.S. dollar ($) and the Canadian dollar (C$) One-Year Forward Exchange Rate Spot Exchange Rate 0.9001 0.8842 Canadian dollar (U.S. dollar/Canadian dollar) The current one-year interest rate on U.S. Treasury securities is 8.03%. If interest rate parity holds, what is the expected yield on one-year Canadian securities of equal risk? 6.12% O 6.43% O5.81% 5.51 %
The current one-year interest rate on U.S. Treasury securities is 8.03 %. If interest rate parity holds, what is the expected yield on one-year Canadian securities of equal risk? O6.12 % 6.43% 5.81% O 5.51% Interest rate parity recognizes that when you invest in a country other than your home country, two factors affect your investment-returns on the investment itself and changes in the exchange rate. Which of the following would cause the overall return on your investment to be lower than the A-Z investment's stated return? Your home currency depreciates relative to the currency in which the investment is denominated. The currency in which the investment is denominated appreciates relative to your home currency. Your home currency appreciates relative to the currency in which the investment is denominated.
Spot Exchange Rate Exchange Rate Canadian dollar (U.S. dolar/Canadian dollar) 0.8842 0.9001 The current one-year interest rate on U.S.. Treasury securities is 8.03 %, If interest rate parity holds, what is the expected yieid on one-year Canadian securities of equal risk? O 6.12% 6.43% 5.81% O 5.51% Interest rate parity recognizes that when you invest in a country other than your home country, two factors affect your investment-returns on the investment itself and changes in the exchange rate. Which of the following would cause the overall return on your investment to be lower than the investment's stated return? Your home currency depreciates relative to the currency in which the investment is denominated. The currency in which the investment is denominated appreciates relative to your home currency Your home currency appreciates relative to the currency in which the investment is denominated
Answer #1

1 CANADIAN $ =0.8442 US $ (SPOT RATE)

1 CANADIAN $ =0.9001 US $ (ONE YEAR FORWARD EXCHANGE RATE)

exchange rate is given in U.S.$ so home country here is U.S. foreign country is Canada

EXPECTED YIELD ON ONE YEAR CANADIAN SECURITIES RISK CAN BE FOUND OUT BY FOLLOWING FORMULA

1 year forward exchange rate / spot exchange rate = (1 + rate in home country) / (1+ expected rate in foreign country)

0.9001/0.8842 = (1+ 0.0803)/(1+ r)

1.018 = 1.0803/1+r

1+r= 1.0803/1.018

r =1.0612-1

=0.0612

=6.12% answer

1.1516 = 1+r

r= .1516